# Quasi Instrumental Variable Methods for Stable Hidden Confounding and Binary Outcomes

### v1

@@ 1 Introduction
In observational studies, identification and estimation of causal effects such as the average treatment effect on the treated (ATT) can be particularly challenging because of concerns about confounding bias. As one can seldom guarantee with certainty that all relevant confounding factors have been measured in a given study, concerns about unmeasured confounding bias has led to several recent developments in causal inference. One such method, the instrumental variable (IV) approach, offers a viable path towards identification of ATT in presence of unmeasured confounding, under key identification conditions; for a recent example, see Liu et al. (2020). Specifically, a valid IV must satisfy three conventional core assumptions: (IV.1) it must be associated with the treatment variable (IV relevance); (IV.2) it must affect the outcome only through the treatment (exclusion restriction); and (IV.3) it must be independent of unmeasured confounders of the treatment-outcome relationship (IV independence). The causal directed acyclic graph (DAG) in Figure (1(a)) provides a graphical depiction of these three core IV assumptions.
@@
Several modes of IV-based causal inference have been proposed over the years for a binary outcome; for examples, see Angrist (2001), Robins and Rotnitzky (2004), Vansteelandt et al. (2011), Clarke and Windmeijer (2012), and Swanson et al. (2018). These methods may be biased if any of core IV assumptions (IV.1)–(IV.3) does not hold. The potential for violation of any of these conditions is a valid concern in practical IV applications; such as so-called Mendelian randomization (MR) studies, a popular form of IV analysis in epidemiology, in which genetic variants (e.g., single-nucleotide polymorphisms or SNPs) are used as instrumental variables to resolve concerns about intractable confounding (Katan, 1986; Davey Smith and Ebrahim, 2003).
@@
A limited but rapidly growing number of methods have recently been developed to address possible violations of assumptions (IV.2) and/or (IV.3) for a continuous outcome in the single IV setting (Spiller et al., 2019, 2022; Tchetgen Tchetgen et al., 2021; Sun et al., 2022; Ye et al., 2024; Liu et al., 2023; Kang et al., 2025). These robust MR methods apply primarily to continuous outcomes, however they do not fully address the unique methodological challenges associated with binary outcomes. First, it is not immediately clear that these methods can uniquely identify the ATT under violations of (IV.2) and/or (IV.3) given that they fundamentally rely on a structural linear model for the outcome. Second, modeling challenges are further accentuated by the presence of nuisance functions which can be difficult to model in a coherent fashion such that they stand a chance for consistent estimation, while ensuring that the resulting ATT estimator lies in its natural parameter space (1,1)(-1,1) for a binary outcome. Third, concerns about potential model mis-specification bias warrants the development of new robust and efficient semiparametric estimators for the causal effect of interest.
@@
To further increase robustness to possible model mis-specification bias, we propose a triply robust semiparametric estimator that is consistent and asymptotically normal (CAN) provided some, but not necessarily all posited nuisance models are correctly specified. We then show that the proposed triply robust estimator attains the semiparametric local efficiency bound when all working models are correctly specified. A further implication of triple robustness in this setting is that it facilitates the use of more flexible machine learning methods via cross-fitting and double/de-biased machine learning (Chernozhukov et al., 2018; Sun et al., 2022; Xu et al., 2022; Cui and Tchetgen Tchetgen, 2023). While we do not pursue these methods in detail in the main text, further details are provided in the Supplemental Material. Finally, we evaluate the finite sample performance of the proposed methods in extensive simulation studies and in an application evaluating the causal effect of adiposity on the risk of hypertension in the UK Biobank.
@@ 2.1 Problem Setup
Suppose that one has observed data O=(Y,A,Z,X)O=(Y,A,Z,{X}), where $A$ is a binary exposure/treatment variable, $Y$ a binary outcome of interest, $Z$ a pre-exposure, possibly invalid binary instrument, and $q$-dimensional baseline covariates ${X}=(X_{1},...,X_{q})$. Let $P(O)$ denote the density of the observed data distribution with respect to the product of the counting measure on $\{0,1\}^{3}$ and $\mu=\otimes_{j=1}^{q}\mu_{j}$, where $\mu_{j}$ is some appropriate dominating measure for the marginal distribution of $X_{j}$. Let $Y_{a}$ denote the potential outcome had possibly contrary to fact, $A$ been set to $a\in\{0,1\}$ and let $Y_{az}$ denote the potential outcome had $A$ and $Z$ been set to $a$ and $z$ respectively, $(a,z)^{T}\in\{0,1\}^{2}$. The causal estimand of interest is the following marginal average treatment effect on the treated (ATT) defined on the additive scale
@@
The conditional expected value of the potential outcome $Y_{a=0}$ among treated individuals $E(Y_{a=0}\mid A=1,{X}={x})$ is not identified because A is potentially subject to unmeasured/hidden confounding even after conditioning on X{X}; this phenomenon is sometimes referred to as the selection problem in the econometrics literature (Angrist and Pischke, 2008), and is also reflected by the confounding bias term being nonzero. In other words, the treatment variable $A$ and the treatment-free potential outcome $Y_{a=0}$ are associated conditional on X{X}, reflecting the presence of confounding bias due to the presence of a hidden common cause of the treatment and outcome.
@@ 2.2 Initial Identification Strategy with an Imperfect IV
An initial strategy to relax (IV.2) and (IV.3) inspired by the above exposition, replaces both conditions with the alternative conditions of stable additive ATT and confounding bias functions with respect to $z$: $\gamma(z,x)=\gamma(x),\delta(z,x)=\delta(x)$; essentially combining the stability conditions of Robins (1994) and Tchetgen Tchetgen and Vansteelandt (2013). This strategy clearly is also relevant to consider in the above DiD setting without imposing the no anticipation assumption. Either way, it is straightforward to check that $E(Y|A=a,Z=z,X=x)$ would then reduce to:
@@ 2.3 An Equilibrium Dynamic Generative Model of Stable Hidden Confounding
The EDG model best captures causal mechanisms in which behaviors and exposures might shift over time due to personal adaptation, environmental pressures, or policy influences only to eventually settle at an equilibrium at which such external and internal forces are at steady state. For example, we can assume that both an individual’s exposure status ($A$) and lifestyle-related confounders (UU)—such as diet, physical activity, or smoking—arise from a dynamic process that changes gradually but whose distribution eventually stabilizes.
@@
To formalize this idea, we now describe a structural equations model. For simplicity, we suppress baseline covariates $X$ and let $p_{aa^{\ast}}(z)=P(Y_{a}=1\mid A=a^{\ast},Z=z)$. Consider the following outcome structural model:
@@
such that $(U,Z)$ is sufficient to account for confounding. While the model clearly accommodates exclusion restriction violations by allowing $\beta_{u,z}\left(U,Z\right)-\beta_{u,z}\left(U,0\right)\neq 0$, it appears to rule out any additive interaction between between $A$ and $Z$ in causing $Y$ conditional on $U$, that is $E(Y_{a=1,z,u}-Y_{a=0,z,u}\mid U=u,Z=z)$ does not depend on $z$. However, we do not view this as a real limitation of the model given that $U$ is not required to be observed and may reasonably be assumed to be enriched with all relevant effect modifiers of the additive effect of $A$ on $Y$, $\beta_{a}\left(u\right)\in(-1,1)$, which is otherwise a priori unrestricted. Furthermore, the no-additive interaction is guaranteed to hold under the sharp null hypothesis of no treatment effect. We also note that in the event $Z$ is a valid instrument, the structural regression model $P\left(Y=1|A=a,U=u\right)=\beta_{a}\left(u\right)a+\beta_{u}\left(u\right)$ is in fact completely unrestricted, other than satisfying the natural constraint that the left hand-side falls within the unit interval $(0,1)$ for all values of $(u,a)$.
@@
does not depend on $u$, where $\beta_{z}\left(0\right)=1$. That is, $\beta_{u,z}\left(u,z\right)=\beta_{u}\left(u\right)\beta_{z}\left(z\right)$. Interestingly, as we show in the Supplementary Materials, this latter assumption may be motivated by a certain conditional independence condition involving counterfactual response types. Finally, we presume that each i.i.d realization $(A,U)$ is a point snapshot of a latent dynamic Markov process, i.e. a unit’s underlying full data actually entails a Markov process $\{A(t),U(t):t\}$ of time-varying treatment subject to time-varying confounding as represented in the DAG given in Figure 2, where, without loss of generality $Z$ is assumed to be realized at $t=0$. The model then assumes that by the time $T$ when the treatment is measured, and thus $A\equiv A(T)\,\$and $U\equiv U(T-1)$ the Markov process has reached a certain equilibrium state, in the sense that its underlying generative law is stable over time, which can formally be stated as the following distributional equality:
@@
such that $\frac{P\left\{Y_{a=0}=1|A=1,z\right\}}{P\left\{Y_{a=0}=1|A=0,z\right\}}$ and $E\left(Y_{a=1}-Y_{a=0}|A=1\right)$ do not depend on $z$. The first condition is a stability condition about the degree of multiplicative confounding as a function that is independent of the QIV, while the second is a stability condition about the degree of effect heterogeneity of the ATT on the additive scale. Both can therefore be seen as implied by the EDG model, although it is important to note that the converse is not necessarily true; i.e. both conditions do not imply the EDG model. As we discuss further in the next Section, the first stability condition can also be interpreted as a familiar nonlinear parallel trends condition with respect to $Z$ (Wooldridge, 2023), in this case a multiplicative parallel trends condition; while the second recovers the well-known no current treatment value interaction assumption of Robins (1994). We further expound upon these conditions below as a basis for credible inference about causal effects in presence of intractable but stable confounding.
@@ 2.4 QIV-based Test of the Causal Null Hypothesis
We propose a novel test of the causal null hypothesis of no conditional ATT,
@@ 2.5 QIV-based Identification of Causal Effects
We have also obtained identification results for causal risk ratio in the treated sub-population, provided in the Supplementary Materials.
@@ 2.6 Lack of Identification with Parallel Trends on the Odds Ratio Scale
Interestingly, in the case of a continuous outcome, Liu et al. (2023) considered the no current treatment value interaction Assumption 3 in conjunction with an alternative parallel trends assumption on the odds ratio scale to establish identification of the additive ATT. In sharp contrast to the case of a continuous outcome, in the Supplementary Materials, we establish the somewhat surprising finding that in case of a binary outcome, the additive ATT may not be uniquely identified if parallel trends is assumed to hold on the odds ratio scale, rather than on the multiplicative scale as in Assumption 2.
@@ 3.1 A Novel Nested Nuisance Model
We obtain the following result that our proposed GOP nuisance model parameterization resolves the variation dependence issue. The proof of Theorem 2 is provided in the Supplementary Materials.
@@
which is contained in the unit interval $(0,1)$. The proof is provided in the supplementary materials. In practice, numerical methods can be employed to find the unique real root within the specified interval (Brent, 2013). Alternatively, the baseline risk $p_{00}(z,x)$ has a closed-form solution derived from the standard theory of cubic equations. We first perform a change of variables following the classical method of Cardano (Abramowitz and Stegun, 1965):
$p_{00}(z,x)=t-\frac{b_{2}(z,x)}{3b_{1}(z,x)},$
which leads to the depressed cubic equation
$t^{3}+\xi\,t+\zeta=0,$
where
$\xi=\frac{3b_{1}(z,x)b_{3}(z,x)-b_{2}(z,x)^{2}}{3b_{1}(z,x)^{2}},\quad\zeta=\frac{2b_{2}(z,x)^{3}-9b_{1}(z,x)b_{2}(z,x)b_{3}(z,x)+27b_{1}(z,x)^{2}b_{4}(z,x)}{27b_{1}(z,x)^{3}}.$
Define the discriminant
$\Delta=\left(\frac{\zeta}{2}\right)^{2}+\left(\frac{\xi}{3}\right)^{3}.$
  • If $\Delta\geq 0$, there is a unique real solution, obtained using Cardano’s formula:
    and
    $p_{00}(z,x)=t-\frac{b_{2}(z,x)}{3b_{1}(z,x)}.$
  • If $\Delta<0$, there are three distinct real roots. Using the trigonometric method (also known as the casus irreducibilis solution):
    $\theta=\cos^{-1}\left(\frac{3\zeta}{2\xi}\sqrt{-\frac{3}{\xi}}\right),$
    the three real solutions are
    $t_{k}=2\sqrt{-\frac{\xi}{3}}\cos\left(\frac{\theta+2k\pi}{3}\right),\quad k=0,1,2,$
    and
    $p_{00}^{(k)}(z,x)=t_{k}-\frac{b_{2}(z,x)}{3b_{1}(z,x)}.$
    We select the root $p_{00}^{(k)}(z,x)$ that falls within the valid interval.
@@ 3.2 Maximum Likelihood Estimation
where $P(Y=1\mid A,Z,X;\varphi)=\tanh\left(\beta\right)A$$\exp\left(\theta A\right)p_{00}\left(Z,{X};\beta,\theta,\omega_{0},\omega_{1},{\boldsymbol{\eta}}\right)$, and the baseline risk $p_{00}\left(Z,{X};\beta,\theta,\omega_{0},\omega_{1},{\eta}\right)$ can be obtained numerically using univariate root-finding algorithm, such as the bisection method (Brent, 2013). The log-likelihood is thus given by
@@
The proposed maximum likelihood estimator has several notable advantages. First, the parametrization is perfectly coherent with model restrictions, so that the model is guaranteed to produce predictions that are within their natural range. Second, the estimator attains the efficiency bound for the posited model in large samples, which is an important benefit, particularly given well known efficiency challenges of MR analyses. Also, because the joint distribution for the treatment and candidate instrument given covariates factorizes, the estimator does not require modeling the latter and is therefore robust to any bias due to misspecification of such a law. Finally, a straightforward score test, or a likelihood ratio test can readily be implemented to test the null hypothesis $H_{0}$ of no conditional ATT. As noted earlier, the approach is guaranteed to have correct type 1 error provided Assumptions 1 and 2 hold, as Assumption 3 holds automatically under $H_{0}$, as long as the likelihood is correctly specified. Although, we expect such a test will be guaranteed to be consistent only against alternatives in a model where Assumption 3 holds, and might have trivial power outside the model. Despite several appealing properties, the MLE does require correct specification of the outcome law, and therefore may be susceptible to bias in case the model is wrong. In the next section we propose a semiparametric estimator with appealing robustness property.

### v2

@@ 1 Introduction
In observational studies, identification and estimation of causal effects such as the average treatment effect on the treated (ATT) can be particularly challenging because of concerns about confounding bias. As one can seldom guarantee that all relevant confounding factors have been measured in a given study, concerns about unmeasured confounding bias have led to several recent developments in causal inference. One such method, the instrumental variable (IV) approach, offers a viable path towards identification of ATT in presence of unmeasured confounding, under key identification conditions; for a recent example, see Liu et al. (2020). Specifically, a valid IV must satisfy three conventional core assumptions: (IV.1) it must be associated with the treatment variable (IV relevance); (IV.2) it must affect the outcome only through the treatment (exclusion restriction); and (IV.3) it must be independent of unmeasured confounders of the treatment-outcome relationship (IV independence). The causal directed acyclic graph (DAG) in Figure (1(a)) provides a graphical depiction of these three core IV assumptions.
@@
Several instrumental variable (IV) methods for causal inference with binary outcomes have been proposed in the literature (Angrist, 2001; Robins and Rotnitzky, 2004; Vansteelandt et al., 2011; Clarke and Windmeijer, 2012; Swanson et al., 2018). These methods may be biased if any of core IV assumptions (IV.1)–(IV.3) does not hold. The potential for violation of any of these conditions is a valid concern in practice; for example, so-called Mendelian randomization (MR) studies, a popular form of IV analysis in epidemiology, in which genetic variants (e.g., single-nucleotide polymorphisms or SNPs) are used as instrumental variables, is susceptible in practice to violations of core IV assumptions (Katan, 1986; Davey Smith and Ebrahim, 2003).
@@
A limited but growing number of methods have been developed for valid inference using an IV despite possible violations of assumptions (IV.2) and/or (IV.3) for a continuous outcome in the single IV setting (Spiller et al., 2019, 2022; Tchetgen Tchetgen et al., 2021; Sun et al., 2022; Ye et al., 2024; Liu et al., 2023; Kang et al., 2025). These robust MR methods apply primarily to continuous outcomes, however they do not fully address the unique methodological challenges associated with binary outcomes. First, it is not immediately clear that these methods can uniquely identify the ATT in analyses of binary outcomes under violations of (IV.2) and/or (IV.3) given that they fundamentally rely on a structural linear model for the outcome. Second, modeling challenges are further accentuated by the presence of nuisance functions which can be difficult to model in a coherent fashion such that they stand a chance for consistent estimation, while ensuring that the resulting ATT estimator lies in its natural parameter space (1,1)(-1,1) for a binary outcome. Third, concerns about potential model misspecification bias warrants the development of new robust and efficient semiparametric estimators for the causal effect of interest.
@@
To further increase robustness to possible model misspecification bias, we propose a triply robust semiparametric estimator that is consistent and asymptotically normal (CAN) provided some, but not necessarily all posited nuisance models are correctly specified. We then show that the proposed triply robust estimator attains the semiparametric local efficiency bound when all working models are correctly specified. Finally, we evaluate the finite sample performance of the proposed methods in extensive simulation studies and in an application evaluating the causal effect of adiposity on the risk of hypertension in the UK Biobank.
@@ 2.1 Problem Setup
Suppose that one has observed data O=(Y,A,Z,X)O=(Y,A,Z,{X}), where $Y$ is a binary outcome of interest, $A$ a binary exposure/treatment variable, $Z$ a pre-exposure, possibly invalid binary instrument, and $q$-dimensional baseline covariates ${X}=(X_{1},...,X_{q})$. Throughout, $U$ will denote a set of unmeasured confounders such that, conditioning on $X$ and $U$ would in principle suffice to identify the causal effect of $A$ on $Y$. Let $P(O)$ denote the density of the observed data distribution with respect to the product of the counting measure on $\{0,1\}^{3}$ and $\mu=\otimes_{j=1}^{q}\mu_{j}$, where $\mu_{j}$ is some appropriate dominating measure for the marginal distribution of $X_{j}$. Let $Y_{a}$ denote the potential outcome had possibly contrary to fact, $A$ been set to $a\in\{0,1\}$ and let $Y_{az}$ denote the potential outcome had $A$ and $Z$ been set to $a$ and $z$ respectively, $(a,z)^{T}\in\{0,1\}^{2}$. The causal estimand of interest is the following marginal average treatment effect on the treated (ATT) defined on the additive scale
@@
The conditional expected value of the potential outcome $Y_{a=0}$ among treated individuals $E(Y_{a=0}\mid A=1,{X}={x})$ is not identified because A is potentially subject to unmeasured/hidden confounding even after conditioning on X{X}; this phenomenon is sometimes referred to as the selection problem in the econometrics literature (Angrist and Pischke, 2008), and is also reflected by the confounding bias term being nonzero. In other words, the treatment variable $A$ and the treatment-free potential outcome $Y_{a=0}$ are associated conditional on X{X}, reflecting the presence of confounding bias due to the presence of $U$, a hidden common cause of the treatment and outcome.
@@ 2.2 Initial Identification Strategy with an Imperfect IV
An initial strategy to relax (IV.2) and (IV.3) inspired by the above exposition, replaces both conditions with the alternative conditions of stable additive ATT and confounding bias functions with respect to $z$: $\gamma(z,x)=\gamma(x),\delta(z,x)=\delta(x)$; essentially combining the stability conditions of Robins (1994) and Tchetgen Tchetgen and Vansteelandt (2013). This strategy clearly is also relevant to consider in the above DiD setting without imposing the no anticipation assumption. Either way, it is straightforward to show that $E(Y|A=a,Z=z,X=x)$ would then reduce to:
@@ 2.3 An Equilibrium Dynamic Generative Model of Stable Hidden Confounding
The EDG model best captures causal mechanisms in which behaviors and exposures might shift over time due to personal adaptation, environmental pressures, or policy influences only to eventually settle at an equilibrium at which such external and internal forces are at steady state. For example, we can assume that both an individual’s exposure status ($A$) and unmeasured lifestyle-related confounders (UU)—such as diet, physical activity, or smoking—arise from a dynamic process that changes gradually but whose distribution eventually stabilizes.
@@
To formalize this idea, we now describe a structural equations model. For simplicity, we suppress baseline covariates $X$ and let $p_{aa^{\ast}}(z)=P(Y_{a}=1\mid A=a^{\ast},Z=z)$. Consider the following outcome structural model that includes the unmeasured confounder of the AAYY association, $U$,:
@@
such that $(U,Z)$ is sufficient to account for confounding. While the model clearly accommodates exclusion restriction violations by allowing $\beta_{u,z}\left(U,Z\right)-\beta_{u,z}\left(U,0\right)\neq 0$, it appears to rule out any additive interaction between $A$ and $Z$ in causing $Y$ conditional on $U$, that is $E(Y_{a=1,z,u}-Y_{a=0,z,u}\mid U=u,Z=z)$ does not depend on $z$. However, we do not view this as a real limitation of the model given that $U$ is not required to be observed and may reasonably be assumed to be enriched with all relevant effect modifiers of the additive effect of $A$ on $Y$, $\beta_{a}\left(u\right)\in(-1,1)$, which is otherwise a priori unrestricted. Furthermore, the no-additive interaction is guaranteed to hold under the sharp null hypothesis of no treatment effect. We also note that in the event $Z$ is a valid instrument, the structural regression model $P\left(Y=1|A=a,U=u\right)=\beta_{a}\left(u\right)a+\beta_{u}\left(u\right)$ is in fact completely unrestricted, other than satisfying the natural constraint that the left hand-side falls within the unit interval $(0,1)$ for all values of $(u,a)$.
@@
does not depend on $u$, where $\beta_{z}\left(0\right)=1$. That is, $\beta_{u,z}\left(u,z\right)=\beta_{u}\left(u\right)\beta_{z}\left(z\right)$. Interestingly, as shown in Section S2 of the Supplementary Materials (“Conditional independence justification for the no-interaction assumption”), this assumption can be motivated by a conditional independence condition involving counterfactual response types.
Finally, we presume that each i.i.d realization $(A,U)$ is a point snapshot of a latent dynamic Markov process, i.e. a unit’s underlying full data actually entails a Markov process $\{A(t),U(t):t\}$ of time-varying treatment subject to time-varying confounding as represented in the DAG given in Figure 2, where, without loss of generality $Z$ is assumed to be realized at $t=0$. The model then assumes that by the time $T$ when the treatment is measured, and thus $A\equiv A(T)\,\$and $U\equiv U(T-1)$ the Markov process has reached a certain equilibrium state, in the sense that its underlying generative law is stable over time, which can formally be stated as the following distributional equality:
@@
such that $\frac{P\left\{Y_{a=0}=1|A=1,z\right\}}{P\left\{Y_{a=0}=1|A=0,z\right\}}$ and $E\left(Y_{a=1}-Y_{a=0}|A=1,z\right)$ do not depend on $z$. The first condition is a stability condition about the degree of multiplicative confounding as a function that is independent of the QIV, while the second is a stability condition about the degree of effect heterogeneity of the ATT on the additive scale. Both can therefore be seen as implied by the EDG model, although it is important to note that the converse is not necessarily true; i.e. both conditions do not imply the EDG model. As we discuss further in the next Section, the first stability condition can also be interpreted as a familiar nonlinear parallel trends condition with respect to $Z$ (Wooldridge, 2023), in this case a multiplicative version of the parallel trends condition of (Tchetgen Tchetgen and Vansteelandt, 2013) and standard DiD (Wooldridge, 2023); while the second recovers the well-known no current treatment value interaction assumption of Robins (1994); both conditions are now directly implied by our Structural EDG model. We further expound upon these conditions below as a basis for credible inference about causal effects in presence of intractable but stable confounding.
@@ 2.4 QIV-based Test of the Causal Null Hypothesis
We propose a novel test of the causal null hypothesis,
@@ 2.5 QIV-based Identification of Causal Effects
We have also obtained identification results for the causal risk ratio in the treated subpopulation, which are presented in Section S1 of the Supplementary Materials.
@@ 2.6 Lack of Identification with Parallel Trends on the Odds Ratio Scale
Interestingly, in the case of a continuous outcome, Liu et al. (2023) considered the no current treatment value interaction Assumption 3 in conjunction with an alternative parallel trends assumption on the odds ratio scale to establish identification of the additive ATT. In sharp contrast to the continuous outcome case, we show in Section S3 of the Supplementary Materials that, for a binary outcome, the additive ATT may not be uniquely identified if parallel trends is assumed to hold on the odds ratio scale rather than on the multiplicative scale as in Assumption 2.
@@ 3.1 A Novel Nested Nuisance Model
We obtain the following result showing that our proposed GOP nuisance model parameterization resolves the variation dependence issue. The proof of Theorem 2 is given in Section S4.2 of the Supplementary Materials.
@@
which is contained in the unit interval $(0,1)$. In practice, numerical methods can be employed to find the unique real root within the specified interval (Brent, 2013). Alternatively, the baseline risk $p_{00}(z,x)$ has a closed-form solution derived from the theory of cubic equations (Abramowitz and Stegun, 1965); the explicit derivation using Cardano’s method is provided in Section S4.3 of the Supplementary Materials.
@@ 3.2 Maximum Likelihood Estimation
where $P(Y=1\mid A,Z,X;\varphi)=\tanh\left(\beta\right)A$$\exp\left(\theta A\right)p_{00}\left(Z,{X};\beta,\theta,\omega_{0},\omega_{1},{\boldsymbol{\eta}}\right)$, and the baseline risk $p_{00}\left(Z,{X};\beta,\theta,\omega_{0},\omega_{1},{\eta}\right)$ can be obtained numerically using a univariate root-finding algorithm, such as the bisection method (Brent, 2013). The log-likelihood is thus given by
@@
The proposed maximum likelihood estimator has several notable advantages. First, the parametrization is perfectly coherent with model restrictions, so that the model is guaranteed to produce predictions that are within their natural range. Second, the estimator attains the efficiency bound for the posited model in large samples, which is an important benefit, particularly given well known efficiency challenges of IV analyses. Also, because the joint distribution for the treatment and candidate instrument given covariates factorizes, the estimator does not require modeling the latter and is therefore robust to any bias due to misspecification of such a law. Finally, a straightforward score test, or a likelihood ratio test can readily be implemented to test the null hypothesis $H_{0}$ of no conditional ATT. As noted earlier, the approach is guaranteed to have correct type 1 error provided Assumptions 1 and 2 hold, as Assumption 3 holds automatically under $H_{0}$, as long as the likelihood is correctly specified. Although, we expect such a test will be guaranteed to be consistent only against alternatives in a model where Assumption 3 holds, and might have trivial power outside the model. Despite several appealing properties, the MLE does require correct specification of the outcome law, and therefore may be susceptible to bias in case the model is wrong. In the next section we propose a semiparametric estimator with appealing robustness property.